# TAPTRv3: Spatial and Temporal Context Foster Robust Tracking of Any Point in Long Video

### v1

@@ Abstract
In this paper, we present TAPTRv3, which is built upon TAPTRv2 to improve its point tracking robustness in long videos. TAPTRv2 is a simple DETR-like framework that can accurately track any point in real-world videos without requiring cost-volume. TAPTRv3 improves TAPTRv2 by addressing its shortage in querying high quality features from long videos, where the target tracking points normally undergo increasing variation over time. In TAPTRv3, we propose to utilize both spatial and temporal context to bring better feature querying along the spatial and temporal dimensions for more robust tracking in long videos. For better spatial feature querying, we present Context-aware Cross-Attention (CCA), which leverages surrounding spatial context to enhance the quality of attention scores when querying image features. For better temporal feature querying, we introduce Visibility-aware Long-Temporal Attention (VLTA) to conduct temporal attention to all past frames while considering their corresponding visibilities, which effectively addresses the feature drifting problem in TAPTRv2 brought by its RNN-like long-temporal modeling. TAPTRv3 surpasses TAPTRv2 by a large margin on most of the challenging datasets and obtains state-of-the-art performance. Even when compared with methods trained with large-scale extra internal data, TAPTRv3 is still competitive.
@@ 1 Introduction
Figure 1: Comparison of TAPTRv3 with prior methods. Evaluation on four challenging test sets, where ‘Avg. L.’ denotes average video length. Circle colors represent training data composition, ‘Real T. Data’ and ‘Syn. T. Data’ denote the real and synthetic data, larger circle indicates more data. Results show TAPTRv3 achieves the best performance on most datasets. Even with only 11K synthetic training data, TAPTRv3 remains competitive against BootsTAPIR that is trained on 15M additional real data.
@@
Localizing points across different frames in a video is a long-standing problem [33]. Recently, with the growing demand for the trajectory and visibility information of arbitrary points in videos for various down-stream tasks, such as video editing [14], SLAM [40], and manipulation [43], the Tracking Any Point (TAP) task has gradually regained attention [12, 7, 20, 26, 25, 30, 46, 45, 41].
@@
To solve this problem, some methods try to construct a 4D field  [45] and track points in the constructed 3D space while also achieving 3D scene perception. Though promising, such methods are normally not general and have inferior performance. By contrast, more methods attempt to solve the TAP task directly from a 2D perspective [12, 7, 20, 26, 25, 41]. Some of these methods follow the traditional optical flow estimation pipeline, RAFT [39] more specifically, and highly rely on cost-volume [48] to perform point tracking like sparse optical flow. Although such methods have achieved impressive performance, the computation of cost-volume is resource-consuming, especially when the number of points, the length of videos, or the video resolution increases.
Figure 2: Illustration of our key contributions. The darker arrow indicates larger attention weight. (a) Point-level feature in TAPTRv2 causes ambiguous attention in cross-attention because of the large variation and distractions. The red point is the tracking point, the green points are sampling points in cross-attention. (b) Patch-level spatial context features help eliminate distractions in TAPTRv3. (c) TAPTRv2 limits temporal attention in a small window. (d) TAPTRv3 extends temporal attention to arbitrary length while utilizing visibility to further adjust the temporal attention weights. Rectangles indicate visibility predictions, the darker the larger.
@@
Inspired by recent visual prompt-based detection [24, 17], TAPTR [26] proposes a DEtection-TRansformer (DETR)-like framework, which regards each tracking point as a point query and addresses the TAP task from the perspective of 2D detection. TAPTRv2 [25] further improves TAPTR by eliminating the cost-volume as it will contaminate the feature of point queries. To compensate for the role of cost-volume for point localization via self-correlation, TAPTRv2 proposes an attention-based position update (APU) operation that utilizes key-aware deformable attention [23] to compare a query point with a set of surrounding sampling points and find a better position. With this improvement, TAPTRv2 leads to both a simpler framework and a better performance.
@@
However, we find that TAPTRv2’s performance in long videos is still not satisfying, due to its shortage of feature querying in both spatial and temporal dimensions in long videos, in which the target tracking points normally undergo increasing variation over time. In the spatial dimension, the key-aware deformation attention used in TAPTRv2 is essentially to compare a query point with a set of surrounding sampling points using point-level features. Such point features are generally too local and often lead to ambiguous matching results due to the existence of repetitive patterns and textureless regions in images (See Fig. 2 (a)). In the temporal dimension, the RNN-like long-temporal modeling in TAPTRv2 often suffers from the drifting problem, as the feature of a tracking point may be gradually affected by ambiguous surrounding features and unknown occlusion over time. Moreover, the current training and testing sets have a large discrepancy in terms of video length [^2] The training set consists of short videos, each having fixed 24 frames, whereas the test set contains videos varying from 50 to 1300 frames.. Excessive feature updates in long videos during testing further exacerbate the feature drifting problem. In our study, we also observe the existence of scene cuts in many videos. While such videos are the result of artificial editing, they are quite prevalent in public datasets  [^3] For instance, in TAP-Vid-Kinetics [6], which is one of the challenging test sets, approximately $27\%$ of the videos contain scene cuts. . The lack of global matching in TAPTRv2 makes it hard to reestablish tracking effectively when a scene cut occurs with sudden large motions.
@@
With these insights, we propose to enhance the feature querying ability of TAPTRv3 in both spatial and temporal dimensions. For spatial feature querying, we develop a Context-aware Cross-Attention (CCA) operation, which replaces the point-level feature comparison employed in key-aware deformable attention with patch-level feature comparison. As illustrated in Fig. 2 (b), such patch-level similarities are more capable of eliminating distractions, making them a more reliable choice for attention weights and leading to a better perception of 2D image space. For temporal feature querying, to address the drifting issue, we discard the RNN-like long-temporal modeling in TAPTRv2. Instead, for any tracking point, we resort to the initial feature sampled from its starting frame where it is specified, as this feature is the most reliable input for point tracking in subsequent frames. To compensate for feature change over time, we introduce a Visibility-aware Long-Temporal Attention (VLTA) operation, which treats the initial feature as a query and performs dense attention over all the past frames to aggregate past feature changes (See Fig. 2 (d)). This not only enables the perception of longer temporal context but also makes TAPTRv3 an online tracker. Meanwhile, recognizing that the target tracking point may be occluded in some frames, we reweight the long-temporal attention weights using the estimated visibility scores in the past frames, directing more attention to frames where the point is visible. This further enhances the feature querying ability along the temporal dimension.
@@
For the scene cut issue, as inspired by previous works [6, 7, 5], we introduce a global matching module to reinitialize the point query’s positional part for subsequent frames. Note that we only trigger the global matching when detecting a scene cut. This is based on our observation that in regular videos, using the predicted positions from the previous frame as the initial position for the current frame yields better results.
@@
In summary, our contributions are threefold: (1) The primary contribution is the development of a more robust solution for point tracking in long videos, which improves upon TAPTRv2 by leveraging both spatial and temporal context. The two corresponding operations, namely Context-aware Cross-Attention and Visibility-aware Long-Temporal Attention, effectively improve local feature matching and long-term feature updating, enhancing feature querying. (2) To address the scene cut issue, we introduce a global matching module, which is only triggered when a scene cut is detected. This ensures stable tracking on regular videos while being able to quickly reestablish tracking when encountering scene cuts. (3) Extensive experimental results show that TAPTRv3 significantly outperforms TAPTRv2 and achieves state-of-the-art performance on most datasets. Even when compared to models trained on large-scale extra internal real data, TAPTRv3 remains competitive.
@@ 2 Related Work
Optical Flow Estimation. Establishing the correspondences of every pixel between two consecutive frames is a long-standing problem. Over the past few decades, extensive research has been dedicated to addressing this issue. Traditional methods [13, 1, 2] use carefully designed descriptors to find correspondences and apply manually designed rules to filter out distractions. DCFlow [48] first demonstrated the feasibility of using the features learned from deep neural network to obtain optical flow estimation through cost-volume that is constructed by calculating 4D correlation, and dominant this field [39, 9, 16, 48, 38, 44, 18, 47, 49, 15, 53]. Although the features extracted by deep neural networks are much stronger, the cost volume still suffers ambiguity. To address this, these methods typically feed the cost-volume into convolutions to normalize it based on contextual information. Although the recent optical flow estimation methods [35, 34, 36] have shown remarkable results, they still can not handle video data well, especially when the interesting points are occluded.
Tracking Any Point. Influenced by optical flow estimation methods, especially the RAFT [39], most approaches [6, 7, 8, 12, 54, 20, 19, 5] follow a similar framework, calculating a cost-volume between the target tracking point and every frame, and then feeding the cost-volume into a transformer [42] to regress the position of the point in each frame. However, the lack of contextual information about the target tracking point, makes the cost-volume a 2D correlation, which is unreliable. Inspired by optical flow methods [39], LocoTrack [5] first points out this limitation, and introduces the local 4D correlation to enhance their performance. In another line, TAPTR [26] and TAPTRv2 [25] address the TAP task from the perspective of detection with their DETR-like [3, 27, 22, 50] framework. With the help of the insight that the cost-volume and attention weights are essentially the same, TAPTRv2 becomes the first method to achieve state-of-the-art performance without requiring the source-consuming cost-volume. However, TAPTRv2 still lacks designs for more challenging long-term tracking, resulting in poor performance in long sequences.
@@ 3 Method
Figure 3: Overview (a) and core components (b) (c) of TAPTRv3. After the user has specified the point to track, the point query preparation stage prepares the content feature and spatial context features for this point in the initial frame. When TAPTRv3 receives a new frame, the sequential point tracking stage will regard the content feature and the specified location as the point query, and regard the new frame’s corresponding image feature map as a set of keys and values. The points query, keys, and values will be sent to the multi-layer transformer decoder to detect the tracking point in the new frame. The location prediction will further be used to update the point query’s positional part, providing a better initial position for tracking in the next frame. For clarity, the global matching module is not plotted.
Since TAPTRv3 shares a similar architecture with TAPTRv2, to make this section self-contained, we will first provide a brief overview of TAPTRv3’s framework and then describe the main improvements of TAPTRv3 in detail.
@@ 3.1 Overview
Before describing TAPTRv3, for clarity and without loss of generality, we assume that only a single point is being tracked, starting from the first frame. Given the first frame of the video $\mathbf{I}_{0}$ and the user-specified point to be tracked at $\mathbf{l}_{0}\in\mathbb{R}^{2}$, TAPTRv3 is expected to detect this point in every subsequent frame $\left\{\mathbf{I}_{t}\right\}_{t=1}^{T-1}$, determining its location $\left\{\mathbf{l}_{t}\right\}_{t=1}^{T-1}$ and visibility $\left\{\alpha_{t}\right\}_{t=1}^{T-1}$, where $\mathbf{l}_{t}\in\mathbb{R}^{2}$, $\alpha_{t}\in[0,1]$, and $T$ is an integer that indicates the length of the video. As shown in Fig. 3 (a), TAPTRv3 can be roughly divided into Point Query Preparation stage and Sequential Point Tracking stage.
Point Query Preparation. As shown in Fig. 3 (a), after the user has specified the point to be tracked on the 00-th frame $\mathbf{I}_{0}$ at $\mathbf{l}_{0}$, the point query preparation stage will sample a point-level feature $\mathbf{f}\in\mathbb{R}^{D}$ to describe the target tracking point, where $D$ is the number of channels. Following TAPTR and TAPTRv2, the sample is achieved by conducting bilinear interpolation on the $\mathbf{I}_{0}$’s corresponding image feature map [^4] For clarity, without loss of generality, we assume that each frame’s feature map, obtained from the backbone and transformer encoder, has only one scale. $\mathbf{X}_{0}\in\mathbb{R}^{H\times W\times D}$ at $\mathbf{l}_{0}$, where $H$ and $W$ indicate the height and width of the feature map. To have a more comprehensive description of the point, TAPTRv3 additionally sample $N^{2}$ context features $\mathbf{C}\in\mathbb{R}^{N^{2}\times D}$ around $\mathbf{l}_{0}$ in a grid form to describe the point’s initial spatial context,
@@
where $N$ is a hyperparameter and is set as 3 by default, $\mathbf{G}\in\mathbb{R}^{N^{2}\times 2}$ is the sampling grid, and Bili is the bilinear interpolation operation. After that, $\mathbf{f}$ and $\mathbf{l}_{0}$ will be regarded as the initial content part and the positional part of the target tracking point’s corresponding point query. The point query will be sent to the transformer decoder as a query, to detect the target tracking point in subsequent frames in the next Sequential Point Tracking stage.

### v2

@@ Abstract
In this paper, built upon TAPTRv2, we present TAPTRv3. TAPTRv2 is a simple yet effective DETR-like point tracking framework that works fine in regular videos but tends to fail in long videos. TAPTRv3 improves TAPTRv2 by addressing its shortcomings in querying high-quality features from long videos, where the target tracking points normally undergo increasing variation over time. In TAPTRv3, we propose to utilize both spatial and temporal context to bring better feature querying along the spatial and temporal dimensions for more robust tracking in long videos. For better spatial feature querying, we identify that off-the-shelf attention mechanisms struggle with point-level tasks and present Context-aware Cross-Attention (CCA). CCA introduces spatial context into the attention mechanism to enhance the quality of attention scores when querying image features. For better temporal feature querying, we introduce Visibility-aware Long-Temporal Attention (VLTA), which conducts temporal attention over past frames while considering their corresponding visibilities. This effectively addresses the feature drifting problem in TAPTRv2 caused by its RNN-like long-term modeling. TAPTRv3 surpasses TAPTRv2 by a large margin on most of the challenging datasets and obtains state-of-the-art performance. Even when compared with methods trained on large-scale extra internal data, TAPTRv3 still demonstrates superiority. Project page: taptr.github.io
@@ 1 Introduction
Localizing points across different frames in a video is a long-standing problem (Sand & Teller, 2008). Recently, with the growing demand for the trajectory and visibility information of arbitrary points in videos for various down-stream tasks, such as video editing (Huang et al., 2023), SLAM (Teufel et al., 2024), and manipulation (Vecerik et al., 2024), the Tracking Any Point (TAP) task has gradually regained attention (Harley et al., 2022; Doersch et al., 2023; Karaev et al., 2024b; Li et al., 2024c; b; Neoral et al., 2024; Xiao et al., 2024; Wang et al., 2023; Tumanyan et al., 2024).
@@
To solve this problem, some methods try to construct a 4D field (Wang et al., 2023) and track points in the constructed 3D space while also achieving 3D scene perception. Though promising, such methods are normally not general and have inferior performance. By contrast, more methods attempt to solve the TAP task directly from a 2D perspective (Harley et al., 2022; Doersch et al., 2023; Karaev et al., 2024b; a; Li et al., 2024c; b; Tumanyan et al., 2024; Aydemir et al., 2025; Zholus et al., 2025). Some of these methods follow the traditional optical flow estimation pipeline, RAFT (Teed & Deng, 2020) more specifically, and highly rely on the dense cost-volume (Xu et al., 2017) to perform point tracking like sparse optical flow. Although such methods have achieved impressive performance, the computation of dense cost-volume is resource-consuming, especially when the number of points, the length of videos, or the video resolution increases.
@@
Inspired by recent visual prompt-based detection methods (Li et al., 2024a; Jiang et al., 2025), TAPTR (Li et al., 2024c) proposes a DEtection-TRansformer (DETR)-like framework, which regards each tracking point as a point query and addresses the TAP task from the perspective of point-level visual prompt detection. TAPTRv2 (Li et al., 2024b) further improves TAPTR by eliminating the requirement of dense cost-volume input, as it will contaminate the point query’s content feature and introduce redundancy. To compensate for its localization role, TAPTRv2 proposes an attention-based position update (APU) operation that utilizes key-aware deformable attention (Li et al., 2023) to compare a query point with a set of key sampling points and find a better position. With this improvement, TAPTRv2 obtains both a simpler framework and a better performance.
@@
However, we find that TAPTRv2 still struggles with long videos due to its shortage of feature querying in both spatial and temporal dimensions in long videos, in which the target tracking points normally undergo increasing variation over time. In the spatial dimension, TAPTRv2 introduces key-aware deformable attention (Li et al., 2023) to extract features and directly perform position update by comparing the feature similarity between a query point and a set of surrounding sampled key points. However, when migrating this module from an object-level detection task to a point-level tracking task, TAPTRv2 overlooks the fact that the point-level query and key features, obtained simply through bilinear interpolation, are too local. This makes the resulting attention weights susceptible to noise. This instability significantly affects point tracking, which requires the most fine-grained spatial understanding. In the temporal dimension, the RNN-like long-temporal modeling in TAPTRv2 often suffers from the drifting problem, as the feature of a tracking point may be gradually affected by ambiguous surrounding features and unknown occlusion over time. Moreover, there is a significant discrepancy in video length between the current training and testing sets. The training set consists of short videos with fixed 24 frames, while the testing videos vary from 50 to 1300 frames in length. Excessive feature updates in long videos during testing further exacerbate the feature drifting problem. In our study, we also observe the existence of scene cuts in many videos. While such videos are the result of artificial editing, they are quite prevalent in public datasets. For instance, in TAP-Vid-Kinetics (Doersch et al., 2022), which is one of the challenging test sets, approximately $27\%$ of the videos contain scene cuts. The lack of global matching in TAPTRv2 makes it hard to reestablish tracking effectively when a scene cut occurs with sudden, large motions.
@@
With these insights, we propose to enhance the feature querying ability of TAPTRv3 in both spatial and temporal dimensions. For spatial feature querying, inspired by the prior 4D cost volume-based optical flow (Teed & Deng, 2020; Ilg et al., 2017; Sun et al., 2018; Wang et al., 2020; Jiang et al., 2021; Xu et al., 2021) and point tracking methods (Cho et al., 2024; Bian et al., 2023), we introduce the context information to the attention mechanism. Specifically, we develop a Context-aware Cross-Attention (CCA) operation to optimize the key-aware deformable attention mechanism. Instead of using unstable point-level similarity, our method leverages patch-level similarity to compute the attention weights. The patch-level attention utilizes more context features to prevent the attention weights from being disturbed, thereby bridging the gap between object-level and point-level tasks. For temporal feature querying, to address the drifting issue, we discard the RNN-like long-temporal modeling in TAPTRv2. Instead, for any tracking point, we resort to the initial feature sampled from its starting frame, as this feature is the most reliable, and use it as input in any frame. To compensate for feature change over time, we introduce a Visibility-aware Long-Temporal Attention (VLTA) operation, which treats the initial feature as a query and performs dense attention over the past frames to aggregate past feature changes. This not only enables the perception of longer temporal context but also makes TAPTRv3 an online tracker. Meanwhile, recognizing that the target tracking point may be occluded in some frames, we reweight the long-temporal attention weights using the estimated visibility scores in the past frames, directing more attention to frames where the point is visible. This further enhances the feature querying ability along the temporal dimension.
@@
For the scene cut issue, we introduce an auto-triggered global matching mechanism to reinitialize the point query’s positional part for subsequent frames. Note that we only trigger the global matching when detecting a scene cut. This is based on our observation that in regular videos, using the predicted positions from the previous frame as the initial position for the current frame yields better results.
@@
In summary, our contributions are threefold: (1) The primary contribution is the development of a more robust solution for point tracking in long videos, which improves upon TAPTRv2 by leveraging both spatial and temporal context. The two corresponding operations, namely Context-aware Cross-Attention and Visibility-aware Long-Temporal Attention, effectively improve the quality of spatial cross attention and long-term feature updating, enhancing feature querying. (2) To address the scene cut issue, we introduce an auto-triggered global matching mechanism, which is only triggered when a scene cut is detected. This ensures stable tracking on regular videos while being able to quickly reestablish tracking when encountering scene cuts. (3) Extensive experimental results show that TAPTRv3 significantly outperforms TAPTRv2 and achieves state-of-the-art performance on most datasets. Even when compared to models trained on large-scale extra internal real data, TAPTRv3 remains competitive.
@@ 2 Related Work
Optical Flow Estimation. Establishing the correspondences for every pixel between two consecutive frames is a long-standing problem. Over the past few decades, extensive research has been dedicated to addressing this issue. Traditional methods (Horn & Schunck, 1981; Black & Anandan, 2002; Bruhn et al., 2005) use carefully designed descriptors to find correspondences and apply manually designed rules to filter out distractions. DCFlow (Xu et al., 2017) first demonstrated the feasibility of using the features learned from deep neural network to obtain optical flow estimation through cost-volume that is constructed by calculating 4D correlation, and dominant this field (Teed & Deng, 2020; Dosovitskiy et al., 2015; Ilg et al., 2017; Xu et al., 2017; Sun et al., 2018; Wang et al., 2020; Jiang et al., 2021; Xu et al., 2021; Zhang et al., 2021; Huang et al., 2022; Zhao et al., 2022). Although the features extracted by deep neural networks are much stronger, the cost-volume still suffers from ambiguity. To address this, these methods typically feed the cost-volume into convolutions to normalize it based on contextual information. Although the recent optical flow estimation methods (Shi et al., 2023a; Saxena et al., 2024; Shi et al., 2023b) have shown remarkable results, they still can not handle video data well, especially when points of interest are occluded.
Tracking Any Point. Influenced by optical flow estimation methods, especially the RAFT (Teed & Deng, 2020), most approaches (Doersch et al., 2022; 2023; 2024; Zheng et al., 2023; Karaev et al., 2024b; a; Cho et al., 2024) follow a similar framework, calculating a cost-volume between the target tracking point and every frame, and then feeding the cost-volume into a transformer (Vaswani et al., 2017) to regress the position of the point in each frame. Inspired by the optical flow method (Teed & Deng, 2020), LocoTrack (Cho et al., 2024) introduces a local 4D correlation to enhance performance. TAG (Harley et al., 2024) extends tracking points to tracking arbitrary targets in videos. AnthroTAP (Kim et al., 2025) proposes a pipeline to generate training labels for point tracking from human motion data. Track-On (Aydemir et al., 2025) focuses on online tracking, introducing memory modules to capture temporal information for reliable point tracking. TAPNext (Zholus et al., 2025) is also an online model that casts this task as sequential masked token decoding and removes tracking-specific inductive biases. In another line, TAPTR and TAPTRv2 (Li et al., 2024c; b) address the TAP task from the perspective of detection with their DETR-like (Carion et al., 2020; Liu et al., 2022; Li et al., 2022; Zhang et al., 2023a) framework. However, TAPTRv2 still lacks designs for more challenging long-term tracking, resulting in suboptimal performance in long sequences.
@@ 3.1 Overview
Before describing TAPTRv3, for clarity and without loss of generality, we assume that only a single point is being tracked, starting from the first frame $\mathbf{I}_{0}$. Given $\mathbf{I}_{0}$ and the user-specified point to be tracked at $\mathbf{l}_{0}\in\mathbb{R}^{2}$, TAPTRv3 is expected to detect this point in every subsequent frame $\left\{\mathbf{I}_{t}\right\}_{t=1}^{T-1}$, determining its location $\left\{\mathbf{l}_{t}\right\}_{t=1}^{T-1}$ and visibility $\left\{\alpha_{t}\right\}_{t=1}^{T-1}$, where $\mathbf{l}_{t}\in\mathbb{R}^{2}$, $\alpha_{t}\in[0,1]$, and $T$ is an integer that indicates the length of the video. As shown in Fig. 1 (a), TAPTRv3 can be roughly divided into the Point Query Preparation stage and the Sequential Point Tracking stage.
Point Query and Spatial Context Preparation. As depicted in Fig. 1 (a), given a user-specified coordinate $\mathbf{l}_{0}$ on the initial frame $\mathbf{I}_{0}$, the point query preparation stage will sample a point-level feature $\mathbf{f}\in\mathbb{R}^{D}$ to describe the target tracking point, where $D$ is the number of channels. Following TAPTRv2, we conduct bilinear interpolation on the $\mathbf{I}_{0}$’s corresponding image feature map [^1] For clarity, without loss of generality, we assume that each frame’s feature map, obtained from the backbone and transformer encoder, has only one scale. $\mathbf{X}_{0}\in\mathbb{R}^{H\times W\times D}$ at $\mathbf{l}_{0}$, where $H$ and $W$ indicate the height and width of the feature map. To create a more comprehensive description of the point, TAPTRv3 additionally samples $N^{2}$ context features $\mathbf{C}\in\mathbb{R}^{N^{2}\times D}$ around $\mathbf{l}_{0}$ in a grid form to describe the point’s initial spatial context:
@@
where $N$ is a hyperparameter and is set as 3 by default, $\mathbf{G}\in\mathbb{R}^{N^{2}\times 2}$ is the sampling grid, and Bili is the bilinear interpolation operation. After that, $\mathbf{f}$ and $\mathbf{l}_{0}$ will be regarded as the initial content part and the positional part of the target tracking point’s corresponding point query. The point query will be sent to the transformer decoder as a query to detect the target tracking point in subsequent frames in the next Sequential Point Tracking stage.
Figure 1: Overview (a) and core components (b) (c) of TAPTRv3. After the user specifies the point to track, the point query preparation stage prepares the content and spatial context features for this point in the initial frame. When TAPTRv3 receives a new frame, the sequential point tracking stage uses the content feature and the specified location as the point query, and regards the new frame’s image feature map as keys and values. The points query, keys, and values are fed into a multi-layer transformer decoder to detect the tracking point in the new frame. The predicted location then updates the point query’s positional part, providing a better initial position for tracking in the next frame. For clarity, the global matching module is not plotted.